OpenStudy (anonymous):

what's the tangent line of (x-1)^2=6(y+2) at point (-5, 4)?

OpenStudy (anonymous):

first i found the derivative... and got (x-1)/3. is that right?

OpenStudy (anonymous):

find dy/dx at (-5,4) it is -2 so slope is -2 y=-2x+k 4=-2*(-5)+k k=-6 y=-2x-6

OpenStudy (anonymous):

yay! kk thanks